OpenStudy (anonymous):

Use mathematical induction to show that a rectangular checkerboard with an even number of cells and two squares missing, one white and one black, can be covered by dominoes.

OpenStudy (anonymous):

this one is best said in english base case two square, \(n=2\) one white one black, can cover with a domino

OpenStudy (anonymous):

now i guess we have to consider the case of an \(m\times n\) rectangular checkerboard where \(m\) is the width and \(n\) is the length. we know that the number of square is even, which means either \(m\) is even or \(n\) is even ( or both) let us say that \(m\) is even remove the last two rows, so now you have an \((m-2)\times n \) checkerboard, which has an even number of tiles, and so by the induction hypothesis, can be covered by dominoes the last two rows is \(2\times n\) and that can clearly be covered by an even number of dominoes either again by induction (so we are using strong induction) or by the obvious fact that you can lay them one across and n up clean this up and it should work

OpenStudy (anonymous):